# Accurate BGV Parameters Selection: Accounting for Secret and Public Key Dependencies in Average-Case Analysis

### v2

@@ 3.4 Ciphertext Addition and Constant Multiplication
and the corresponding critical quantity scales by $k$, yielding
$\nu_{\mathsf{const}}=k\cdot\nu.$
In both cases, decryption correctness is preserved as long as the resulting noise does not cause the message to wrap around the modulus.
3.5 Homomorphic multiplication and Key switching
Homomorphic multiplication is one of the fundamental operations supported by the BGV scheme. Given two ciphertexts $\bm{c}=(c_{0},c_{1})$ and $\bm{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime})$, and assuming that both are defined with respect to the same modulus $q_{l}$, their product is given by
$\mathsf{Mul}(\bm{c},\bm{c^{\prime}}):=(c_{0}^{\mathsf{mul}},c_{1}^{\mathsf{mul}},c_{2}^{\mathsf{mul}})=(c_{0}\cdot c_{0}^{\prime},c_{0}\cdot c_{1}^{\prime}+c_{1}\cdot c_{0}^{\prime},c_{1}\cdot c_{1}^{\prime})\mod q_{l}.$
@@ 3.5 Homomorphic multiplication and Key switching
As a result, the ciphertext expands from two to three polynomials, which violates the compactness property and makes subsequent operations more costly. This phenomenon becomes even more problematic when performing multiple multiplications. Recovering the message from $\mathsf{Mul}(\bm{c},\bm{c^{\prime}})$ requires computing the reduction modulo $t$ of the resulting critical quantity given by $\nu_{\mathsf{mul}}=c_{0}^{\mathsf{mul}}+c_{1}^{\mathsf{mul}}s+c_{2}^{\mathsf{mul}}s^{2}$.
It is significant to note that, unlike other operations, multiplication causes the noise to grow multiplicatively, making it the most critical case to handle among the basic homomorphic operations in terms of error growth.
To address this, a technique known as relinearization, or key switching, is employed. The key idea is to reduce the ciphertext back to its original size while preserving its correctness. In practice, this means converting a ciphertext of the form $(c_{0},c_{1},c_{2})$ into a new, equivalent one of the form $(\hat{c}_{0},\hat{c}_{1})$, using auxiliary data called the key switching key. More intuitively, key switching allows the expression $c_{0}+c_{1}s+c_{2}s^{2}$ to be re-expressed as a linear polynomial in $s$, restoring the standard noise form.
@@
Although the relinearization step introduces additional noise, it is essential for maintaining the practicality of the scheme. Several key switching techniques have been proposed in the literature, each aiming to optimize this trade-off between correctness and noise growth. The most commonly used are the Brakerski Vaikuntanathan (BV) variant, the Gentry Halevi Smart (GHS) variant and the Hybrid variant, which can be considered as a combination of the previous ones.
3.6 Modulus switching
The primary aim of the modulus switching technique is to reduce the noise resulting from homomorphic evaluations. In practical circuits, modulus switching is typically applied only after homomorphic multiplications, due to their higher cost in terms of error growth. Similarly to the bootstrapping technique, the main purpose is to transition from an extended ciphertext $(\bm{c},q_{l},\nu)$ to another $(\bm{c}^{\prime},q_{l^{\prime}},\nu^{\prime})$ such that the error associated to the latter is smaller, while preserving the fact that, using the same key, both can be decrypted to the same original message. The key difference lies in how this is achieved, and it is in this aspect that modulus switching demonstrates greater efficiency compared to bootstrapping.
@@ 3.6 Modulus switching
Let $(\bm{c},q_{l},\nu)$ be the extended ciphertext, whose error we want to reduce, and let $l^{\prime}$ be an integer, such that $q_{l^{\prime}}<q_{l}$. The new ciphertext produced by the modulus switch is given by
@@
where $\bm{\delta}=t[-\bm{c}t^{-1}]_{\frac{q_{l}}{q_{l^{\prime}}}}$. The $\bm{\delta}$ value can be interpreted as a correction required to ensure that the ciphertext is divisible by $\frac{q_{l}}{q_{l^{\prime}}}$ and does not affect the original message. In fact, it only influences the error since $\bm{\delta}\equiv 0\mod t$. Therefore, the new ciphertext $\bm{c}^{\prime}$ will still decrypt to the original plaintext (scaled by a factor of $\frac{q_{l}}{q_{l^{\prime}}}$), provided that the necessary conditions are met.
@@
The critical quantity associated to the new ciphertext $\bm{c^{\prime}}$ can be expressed in terms of that of $\bm{c}$, namely $\nu$, as
$\nu_{\mathsf{modswitch}}$ $=[c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{l^{\prime}}}=\frac{q_{l^{\prime}}}{q_{l}}([c_{0}+c_{1}\cdot s]_{q_{l}}+\delta_{0}+\delta_{1}\cdot s)=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\delta_{0}+\delta_{1}\cdot s).$
To conclude, it is worth noting that this correctness verification implicitly relies on the common value $h$ in the two equalities
$\begin{split}[c_{0}+c_{1}\cdot s]_{q_{l}}\hskip 2.84544pt&=c_{0}+c_{1}\cdot s-hq_{l}\\ [c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{l^{\prime}}}&=c_{0}^{\prime}+c_{1}^{\prime}\cdot s-hq_{l^{\prime}}\end{split}$
The reason why the value $h$ is the same for both the expressions can be found in [4, Lemma 1].
@@ 4 Average-Case Noise Analysis for BGV
The aim of this section is to investigate the behavior of the noise resulting from the main homomorphic operations supported by the BGV scheme. Throughout this analysis, we consider ciphertexts that are mutually independent, obtained by encrypting independently generated random messages under the same public key.
As previously mentioned, the novel approach introduced in this paper seeks to analyze noise growth by accounting for dependencies among the coefficients of the critical quantities involved. Before delving into the details, it is important to highlight that in BGV, the critical quantity resulting from homomorphic operations can be affected by such dependencies, even when the ciphertexts involved are independent. These dependencies stem from the fact that the noise in the ciphertexts includes terms involving powers of the secret key $s$ and powers of the error term $e$, which makes it necessary to explicitly consider these contributions when analyzing the variance of the noise.
@@
To enhance clarity, for the critical quantity $\nu$ of a fresh ciphertext $\bm{c}$, this notation yields
$\nu=a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s,$
@@
From this point onward, we adopt the notation $n$ to denote $\phi(m)$, in order to simplify the presentation. It is interesting to point out that when $n$ is a power of two, $m=2n$.
4.1 Distribution Analysis
As previously mentioned, the noise coefficients are treated as random variables. We observed that their distributions are well-approximated by identically distributed, dependent Gaussian variables centered at zero. This empirical observation was validated through experiments using the Python fitter package, which confirmed the Gaussian nature of the noise coefficients. This property is particularly advantageous, as it allows us to bound the maximum absolute value of the noise coefficients with high probability by simply controlling their variance $V$. In particular, we recall the following fundamental property of Gaussian random variables.
Proposition 3
Let $Z$ be a real-valued, zero-mean Gaussian random variable with variance $V$. Then, for any fixed value $z>0$, the probability that $Z$ lies within the interval $(-z,z)$ is given by
$\mathbb{P}(|Z|<z)=\text{erf}\left(\frac{z}{\sqrt{2V}}\right),$
where $\text{erf}(z)$ is the error function, defined as
[…]
Extending this result to a vector $\bm{Z}\in\mathbb{R}^{n}$, whose entries are distributed as $Z$, we obtain the following bound on the tail probability of the infinity norm
$\mathbb{P}(\|\bm{Z}\|_{\infty}>z)\leq n\left(1-\text{erf}\left(\frac{z}{\sqrt{2V}}\right)\right).$ (2)
This implies that, in order to ensure that a noise vector $\nu$ satisfies the correctness condition $\|\nu\|_{\infty}<{q}/{2},$ we can use Proposition 3 to bound the failure probability as
where $V$ denotes the estimated variance of the noise coefficients. To express this bound more conveniently, we introduce a security parameter $D$ such that
$D\leq\frac{q}{2\sqrt{2V}},$
from which it follows, using the monotonicity of the error function, that
Consequently, the ciphertext modulus $q$ can be selected according to the inequality
$q\geq 2D\sqrt{2V},$ (3)
which highlights the importance of accurately estimating $V$.
It is important to emphasize that the bounds derived in our analysis provide insight into the minimum ciphertext modulus $q$ required to guarantee the correctness of the scheme. Although larger values of $q$ may be used in practice, identifying the minimal admissible value is essential for optimizing performance and efficiency. This is particularly relevant in the context of the BGV scheme, where our results can serve as a guideline for carefully selecting the setting parameters. By determining tight lower bounds on $q$, it is possible to lay the foundation for sound and efficient parameter selection that ensures correctness without over-provisioning resources.
@@ 4.2 Mean and Variance Analysis
As previously discussed, we have demonstrated that the coefficients of the error term are centered at zero. Furthermore, we have shown that the coefficients of the $b_{\mu}$ terms are uncorrelated, meaning that their pairwise covariance is zero.
@@
Let $\nu=\sum_{\iota}a_{\iota}s^{\iota}$ represent the critical quantity associated with a given ciphertext, where, for a fixed $\iota$, $a_{\iota}=\sum_{\mu}b_{\mu}(\iota)e^{\mu}$. Then, the following equivalence holds
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}]$
@@ 4.3 Homomorphic multiplications
The main goal of our analysis is to provide estimates on the growth of the error resulting from homomorphic multiplications, which are the most complex operations and therefore the most interesting to study. It thus becomes crucial to account for the dependencies that arise among the noise coefficients as a consequence of these multiplications. To this end, the approach we adopted relies on the correction functions introduced in [2].
Therefore, before stating the theorem, some of the key properties of $F$ will be recalled. Unlike the BFV scheme, the BGV scheme requires additional care in the use of the correction function $F$, as not only the impact of the term $s$ must be taken into account, but also that of the term $e$, which is negligible in BFV. Consequently, our analysis requires the introduction of two distinct correction functions, which will be denoted as $F_{e}$ and $F_{s}$. To define these two functions, we need the following
Heuristic 1
Let $x$ be a polynomial whose coefficients are independent and identically distributed according to a fixed distribution. The function $f_{x}$ defined as
$f_{x}(\iota)=-e^{\alpha-\beta\iota-\gamma\iota^{2}}+\delta,$ (4)
where $\alpha,\beta,\gamma,\delta$ depend only on the coefficient distribution of $x$ and the ring dimension $n$, satisfies the following
$f_{x}(\iota)\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[x^{\iota}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[x^{\iota-1}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[x|_{i_{2}}^{2}]}.$
In Table 1, we report the values of $\alpha,\beta,\gamma,\delta$ required to define the function $f_{s}$, where $s$ denotes the secret key with coefficients sampled from the ternary distribution $\chi_{s}=\mathcal{U}_{3}$, as in [2].
        $n$         $\alpha$         $\beta$         $\gamma$         $\delta$
        $2^{12}$         2.8732         0.0160         0.0049         19.1895
        $2^{13}$         2.9644         0.0196         0.0046         20.4747
        $2^{14}$         2.9578         0.0386         0.0032         19.5755
        $2^{15}$         2.9765         0.0197         0.0043         20.7760
Table 1: Parameters for $\chi_{s}=\mathcal{U}_{3}$
Similarly, it is possible to define the function $f_{e}$ for the error term $e$, whose coefficients are distributed according to a discrete Gaussian distribution. For completeness, a table containing the parameters depending on the distribution $\chi_{e}$ can be found in Table 2 [2].
        $n$         $\alpha$         $\beta$         $\gamma$         $\delta$
        $2^{12}$         2.9000         0.0157         0.0051         19.5356
        $2^{13}$         2.9340         0.0042         0.0055         20.7063
        $2^{14}$         2.9138         0.0290         0.0039         19.2973
        $2^{15}$         2.9511         0.0129         0.0046         20.7263
Table 2: Parameters for $\chi_{e}=\mathcal{DG}_{q}(\sigma^{2})$, $\sigma=3.19$
@@
Let $g_{x}(\iota)=\prod_{i=0}^{\iota}f_{x}(i)$ with $f_{x}$ defined as in (4) and $f_{x}(0)=f_{x}(1)=~{}1$. The correction function $F_{x}$ is
$F_{x}(\iota_{1},\iota_{2}):=\frac{g_{x}(\iota_{1}+\iota_{2})}{g_{x}(\iota_{1})\cdot g_{x}(\iota_{2})}$
Moreover, this function satisfies
$F_{x}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[x^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[x^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[x^{\iota_{2}}|_{i_{2}}^{2}]}$
A proof of this result can be found in [2]. For the purposes of our study, we denote by $F_{e}$ the correction function required to account for the dependencies induced by the term $e$, and by $F_{s}$ the correction function associated with the term $s$.
@@
where $K_{1},K_{2}$ arise from the expansions
$a_{\iota_{1}}=\sum_{\mu_{1}=0}^{K_{1}}b_{\mu_{1}}(\iota_{1})e^{\mu_{1}},\hskip 7.11317pta^{\prime}_{\iota_{2}}=\sum_{\mu_{2}=0}^{K_{2}}b^{\prime}_{\mu_{2}}(\iota_{2})e^{\mu_{2}}$
and represent the highest power of $e$ appearing in $a_{\iota_{1}},a^{\prime}_{\iota_{2}}$, respectively.
@@
$n\left(\sum_{\mu_{1}=0}^{K_{1}}\mathsf{Var}(b_{\mu_{1}}(\iota_{1})|_{i})\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{2}}]\right)\cdot$
$\cdot\left(\sum_{\mu_{2}=0}^{K_{2}}\mathsf{Var}(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i})\sum_{j_{3}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}]\right),$
which can be reordered as
@@
Recalling that the correction function $F$ approximates the following quantities
$\begin{cases}F_{e}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[e^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[e^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[e^{\iota_{2}}|_{i_{2}}^{2}]}\\ F_{s}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[s^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|_{i_{2}}^{2}]}\par\par\end{cases}$
It is possible to use these functions in order to express $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ in terms of $\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$. In fact, $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ can be written as

### v3

@@ 3.2 Homomorphic Operations
and the corresponding critical quantity scales by $k$, yielding $\nu_{\mathsf{const}}=k\cdot\nu$.
Homomorphic multiplication.
Let $\bm{c},\bm{c^{\prime}}$ be two ciphertexts defined in $\mathcal{R}_{q_{\ell}}$, then
$\mathsf{Mul}(\bm{c},\bm{c^{\prime}}):=(d_{0},d_{1},d_{2})=(c_{0}\cdot c_{0}^{\prime},c_{0}\cdot c_{1}^{\prime}+c_{1}\cdot c_{0}^{\prime},c_{1}\cdot c_{1}^{\prime})\mod q_{\ell}.$
@@
As a result, the ciphertext expands from two to three polynomials, which violates the compactness property and makes subsequent operations more costly.
Recovering the message from $\mathsf{Mul}(\bm{c},\bm{c^{\prime}})$ requires computing the reduction modulo $t$ of the resulting critical quantity given by $\nu_{\mathsf{mul}}=d_{0}+d_{1}s+d_{2}s^{2}$.
To modify the ciphertext polynomial $d_{0}+d_{1}s+d_{2}s^{2}$ back to another polynomial $\bar{c}_{0}+\bar{c}_{1}\cdot s$ encrypting the same plaintext, a technique known as relinearization, or key switching, is employed.
Key switching.
Intuitively, the key switching technique converts $d_{2}s^{2}$ into $\hat{c}_{0}+\hat{c}_{1}s$ using somehow the encryption of $s^{2}$ under $s$. Indeed,
$\mathsf{Enc}_{s}(s^{2})=(\beta,\alpha)=(-us+te+s^{2},u+te_{1})\approx(\alpha s+s^{2},-\alpha).$
Thus, $s^{2}\approx\beta-\alpha s$ and then $d_{0}+d_{1}s+d_{2}s^{2}\approx d_{0}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+d_{1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s+d_{2}(\beta-\alpha s)=\bar{c}_{0}+\bar{c}_{1}s$. As expected, while the relinearization step introduces additional noise, it is crucial for ensuring the practicality of the scheme.
@@
Several key switching techniques have been proposed in the literature, each aiming to optimize this trade-off between correctness and noise growth. The most commonly used are the Brakerski Vaikuntanathan (BV) variant [7], the Gentry Halevi Smart (GHS) variant [25], and the Hybrid variant [25], which can be considered as a combination of the previous ones. Herein, we do not delve into the details of each method and refer the reader to [34] for further information.
Modulus switching.
The primary aim of the modulus switching technique is to reduce the noise resulting from homomorphic operations.
@@
Let $(\bm{c},q_{\ell},\nu)$ be the extended ciphertext whose error we aim to reduce, and let $\ell^{\prime}$ be an integer such that $q_{\ell^{\prime}}<q_{\ell}$. The modulus switching procedure outputs $(\bm{c}^{\prime},q_{\ell^{\prime}},\nu^{\prime})$, where
$\bm{c}^{\prime}=\frac{q_{\ell^{\prime}}}{q_{\ell}}(\bm{c}+\bm{\delta})\mod q_{\ell^{\prime}},$
@@
with $\bm{\delta}=t[-\bm{c}t^{-1}]_{{q_{\ell}}/{q_{\ell^{\prime}}}}$. The $\bm{\delta}$ value can be interpreted as a correction required to ensure that the ciphertext is divisible by ${q_{\ell}}/q_{\ell^{\prime}}$ and does not affect the original message. In fact, it only influences the error since $\bm{\delta}\equiv 0\mod t$. Therefore, the new ciphertext $\bm{c}^{\prime}$ will still decrypt to the original plaintext (scaled by a factor of ${q_{\ell}}/q_{\ell^{\prime}}$).
@@
The critical quantity associated to the new ciphertext $\bm{c^{\prime}}$ can be expressed in terms of that of $\bm{c}$, namely,
$\nu_{\mathsf{ms}}$ $=[c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{\ell^{\prime}}}=\frac{q_{\ell^{\prime}}}{q_{\ell}}([c_{0}+c_{1}\cdot s]_{q_{\ell}}+\delta_{0}+\delta_{1}\cdot s)=\frac{q_{\ell^{\prime}}}{q_{\ell}}(\nu+\delta_{0}+\delta_{1}\cdot s).$
@@ 4 Average-Case Noise Analysis for BGV
The aim of this section is to investigate the behavior of the noise resulting from the main homomorphic operations supported by the BGV scheme. Throughout this analysis, we consider ciphertexts that are mutually independent, obtained by encrypting independently generated random messages using the same public key.
As previously mentioned, the novel approach introduced in this paper seeks to analyze noise growth by accounting for dependencies among the coefficients of the critical quantities involved. Before delving into the details, it is important to highlight that in BGV, the critical quantity resulting from homomorphic operations can be affected by such dependencies, even when the ciphertexts involved are independent. These dependencies arise because the noise in the ciphertexts includes terms involving powers of the secret key $s$ and powers of the error term $e$, which makes it necessary to explicitly consider these contributions when analyzing the variance of the noise.
@@
To enhance clarity, for the critical quantity $\nu_{\mathsf{clean}}$ of a fresh ciphertext $\bm{c}$, this notation yields
$\nu_{\mathsf{clean}}=a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s,$
@@ 4.1 Mean and Variance Analysis
In the following, we prove that the coefficients of the error term are centered at zero. Furthermore, we show that the coefficients of the $b_{\mu}$ terms are uncorrelated, meaning that their pairwise covariance is zero.
@@
Let $\nu=\sum_{\iota}a_{\iota}s^{\iota}$ represent the critical quantity associated with a given ciphertext, where, for a fixed $\iota$, $a_{\iota}=\sum_{\mu}b_{\mu}(\iota)e^{\mu}$. Then, the following identity holds
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}].$
Moreover, $\mathsf{Var}(\nu|_{i})=\sum_{\iota,\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}].$
@@
4.2 Homomorphic Operations
We can state our results on the variance computation for operations, especially focusing on the multiplication in the next section.
Proposition 1(Encryption)
The invariant noise $\nu_{\mathsf{clean}}$ of a fresh ciphertext has coefficient variance
$V_{\mathsf{clean}}=\mathsf{Var}(\nu_{\mathsf{clean}}|_{i})=\tfrac{t^{2}}{q^{2}}\left(\tfrac{1}{12}+nV_{e}V_{u}+V_{e}+nV_{e}V_{s}\right).$ (4)
Proof
The fresh error $\nu_{\mathsf{clean}}$ can be written as $a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s$, where
$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}.\end{cases}$
The proof is thus concluded by applying Lemma 2 and observing that $\mathbb{E}[s|_{j}^{2}]=\mathsf{Var}(s|_{j})=V_{s}$ and $\mathbb{E}[e|_{j}^{2}]=\mathsf{Var}(e|_{j})=V_{e}$ since $\mathbb{E}[s|_{j}]=\mathbb{E}[e|_{j}]=0$.
Proposition 2(Addition & Constant Multiplication)
Let $\alpha\in\mathcal{R}_{t}$ and $\mathbf{c}$, $\mathbf{c}^{\prime}$ be two independently-computed ciphertexts with invariant noises $\nu,$ and $\nu^{\prime}$, respectively. Then, the variance of the error coefficients
  • resulting from the addition of $\mathbf{c}$ and $\mathbf{c}^{\prime}$ is
    $\mathsf{Var}(\nu_{\mathsf{add}}|_{i})=\mathsf{Var}(\nu|_{i})+\mathsf{Var}(\nu^{\prime}|_{i}).$ (5)
  • after a multiplication between $\alpha$ and $\mathbf{c}$ is
    $\mathsf{Var}(\nu_{\mathsf{const}}|_{i})=\tfrac{(t^{2}-1)n}{12}\mathsf{Var}(\nu|_{i}).$ (6)
Proposition 3(Modulo Switch)
Let $\mathfrak{c}=(\mathbf{c},q_{\ell},\nu)$ be an extended ciphertext. The variance of the error coefficients after the modulo switch to the target modulo $q^{\prime}_{\ell}$ is
$V(\nu_{\mathsf{ms}}|_{i})=\tfrac{q_{\ell}^{\prime 2}}{q_{\ell}^{2}}\mathsf{Var}(\nu|_{i})+\tfrac{t^{2}}{12}(1+nV_{s}).$ (7)
Propositions 2 and 3 follow directly from Lemma 2 by arguments analogous to those in Proposition 1. The proofs are therefore omitted.
We do not explicitly address key switching here, since its negligible impact is discussed in Appendix 0.C.
@@ 4.3 Homomorphic multiplications
The main goal of our analysis is to estimate the error growth caused by homomorphic multiplications, the most critical operations to handle. This requires accounting for the dependencies among noise coefficients induced by such multiplications. Our method relies on a correction function $F$, whose construction is detailed in this section. We recall that the function $F$ was introduced in [4] to “correct” the product of the variances, since these are not independent. Unlike BFV, the BGV scheme demands particular care, as both terms $s$ and $e$ must be considered, whereas the contribution of $e$ is negligible in BFV. Furthermore, unlike [4], our construction of $F$ is not heuristic but is based on Lemma 3. We recall the Isserlis theorem which will be exploited in the proof of the Lemma.
Theorem 4.1
[32, Chapter 8] Let $(X_{1},\ldots,X_{n})$ be a zero-mean multivariate normal random vector, then
$\mathbb{E}[X_{1}\cdots X_{n}]=\sum\prod_{i,j}\mathbb{E}[X_{i}X_{j}]\,$,
where the sum is over all the partition of $\{1,\ldots,n\}$ into pairs and the $(i,j)$ ranges in these pairs.
Lemma 3
Let $a(x)=a|_{0}+a|_{1}x+\dots+a|_{n-1}x^{n-1}\in\mathcal{R}$, where $a|_{i}$’s are i.i.d. random variables with $\mathbb{E}[a|_{i}]=0$, $\mathbb{E}[a|_{i}^{2}]=V_{a}$.
Then
$\mathbb{E}[a^{k}|_{i}]=0,\quad\mathbb{E}[(a^{k}|_{i})^{2}]=k!n^{k-1}V_{a}^{k}\,$,
for all $0\leq i\leq n-1$, where $a^{k}|_{i}$ denotes the ii–th coefficient of the polynomial $a(x)^{k}$, for $n$ sufficiently large.
Proof
Without loss of generality, we prove the statement for $a^{k}|_{0}$, since the same argument applies to all other terms. Let $\omega_{1},\dots,\omega_{n}$ be the roots of $x^{n}+1$, which is a cyclotomic polynomial since $n$ is a power of two. Then, we know that $\sum_{i=1}^{n}\omega_{i}^{m}=0$ for $1\leq m\leq n-1$, and $a(\omega_{1})^{k}+\dots+a(\omega_{n})^{k}=na^{k}|_{0}+a^{k}|_{1}(\omega_{1}+\dots+\omega_{n})+\dots+a^{k}|_{n-1}(\omega_{1}^{n-1}+\dots+\omega_{n}^{n-1})$, thus
$a^{k}|_{0}=\frac{1}{n}(a(\omega_{1})^{k}+\dots+a(\omega_{n})^{k}).$
Now, we define $Z_{i}:=a(\omega_{i})=a_{0}+a_{1}\omega_{i}+\dots+a_{n-1}\omega_{i}^{n-1},$ for $i=1,\ldots,n,$ where, for $n$ sufficiently large, thanks to [20, Theorem 9] we have that $Z_{i}$ follows a Gaussian distribution centered at zero. Then
$\mathbb{E}[Z_{m}\cdot Z_{l}]$ $=\mathbb{E}[(a|_{0}+a|_{1}\omega_{m}+\dots+a|_{n-1}\omega_{m}^{n-1})(a|_{0}+a|_{1}\omega_{l}+\dots+a|_{n-1}\omega_{l}^{n-1})]=$
$=\mathbb{E}[a|_{0}^{2}]+\mathbb{E}[a|_{1}^{2}]\omega_{m}\omega_{l}+\dots+\mathbb{E}[a|_{n-1}^{2}]\omega_{m}^{n-1}\omega_{l}^{n-1}=$
$=V_{a}(1+\omega_{m}\omega_{l}+\dots+\omega_{m}^{n-1}\omega_{l}^{n-1})=$
$=\begin{cases}nV_{a}\quad\quad\text{if}\quad\omega_{m}=\omega_{l}^{-1}\\ 0\quad\quad\text{otherwise}.\end{cases}$
Therefore, for $m=l$, $\mathbb{E}[Z_{m}^{2}]=0$ and, since every linear combination of $Z_{i}$’s is normally distributed, we can apply the Isserlis theorem (Theorem 4.1) to obtain $\mathbb{E}[Z_{m}^{k}]=0$. Consequently
$\mathbb{E}[a^{k}|_{0}]=\frac{1}{n}(\mathbb{E}[Z_{1}^{k}]+\dots+\mathbb{E}[Z_{n}^{k}])=0.$
Moreover, $\mathbb{E}[a^{k}|_{0}^{2}]=\frac{1}{n^{2}}\sum_{m}\mathbb{E}[Z_{m}^{k}Z_{l}^{k}]=\frac{1}{n^{2}}\sum_{m,l}\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]=k!n^{k-1}V_{a}^{k}$, where the last equality follows again from the Isserlis theorem. Indeed,
$\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]$ $=\mathbb{E}[Z_{m}\cdot\dots\cdot Z_{m}\cdot Z_{m^{-1}}\cdot\dots\cdot Z_{m^{-1}}]=$
$=\mathbb{E}[X_{1}\cdot\dots\cdot X_{m}\cdot X_{m+1}\cdot\dots\cdot X_{2m}]=$
$=\sum\prod\mathbb{E}[X_{i}X_{j}]=$
$=\begin{cases}0\quad\quad\text{if}\,\,\,0\leq i,j\leq k\,\,\,\text{or}\,\,\,k+1\leq i,j\leq 2k+1\\ nV_{a}\quad\quad\text{otherwise}.\end{cases}$
Therefore, $\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]$ is given by the sum of $k!$ addends whose value is $n^{k}V_{a}^{k}$, i.e., $\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]=k!n^{k}V_{a}^{k}$. Finally, $\mathbb{E}[a^{k}|_{0}^{2}]=\frac{1}{n^{2}}\cdot n\cdot k!\cdot n^{k}\cdot V_{a}^{k}=k!n^{k-1}V_{a}^{k}$. ∎
Remark 1
This result extends some of [24, Theorem 4.3], where the authors proved the same statement, using a different approach, but only for the case where the $a|_{i}$’s are normal independent random variables. In our case, the coefficients of the polynomial $a(x)$ follow any distribution (with zero mean and finite variance). Moreover, a proof of the specific case $k=2$ is also provided in [5, Lemma 2].
@@
Let $a\in\mathcal{R}$. The correction function $F_{a}$ is defined as
$F_{a}(\iota_{1},\iota_{2})=\frac{\sum_{i=0}^{n-1}\mathbb{E}[a^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[a^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[a^{\iota_{2}}|_{i_{2}}^{2}]}\,.$ (8)
From Lemma 3, it follows that for polynomials with coefficients following distributions of that form, it holds that
$F_{a}(\iota_{1},\iota_{2})=\frac{(\iota_{1}+\iota_{2})!}{\iota_{1}!\,\iota_{2}!}\,.$ (9)
We observe that the value of $F_{a}$ is the same for all polynomials as in Lemma 3. However, for the sake of clarity in presenting our results, we will keep separate notations, writing $F_{s}$ for the secret key $s$ and $F_{e}$ for the public key $e$.
@@
By (8), we express $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ in terms of $\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$. In fact, $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ can be written as